OpenStudy (anonymous):

The 9 officers of the student council are going on an tripan amusement park.Each must pay an entrance fee plus $5for meals. The total cost of the trip is $225.Sovle the equation 9(e+5)=225 to find the cost of the entrance fee for each student.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

9e + 45 = 225 9e = 225 - 45 9e = 180 e = 180/9 e = 20